A bottle of perfume sells for $84.97 and costs the merchant $25. What is the markup rate?
Answer by fractalier(2101)

Selling price = cost + markup
But markup = markup rate * cost, thus
Selling price = cost + (markup rate * cost)
Now plug in
84.97 = 25 + 25r
59.97 = 25r
r = 59.97 / 25 or almost 2.4, which is 240%